### v2

# Coloring Hardness on Low Twin-Width Graphs

Abstract
As the class $\mathcal{T}_{4}$ of graphs of twin-width at most 4 contains every finite subgraph of the infinite grid and every graph obtained by subdividing each edge of an $n$-vertex graph at least $2\log n$ times, most $\mathsf{NP}$-hard graph problems, like Max Independent Set, Dominating Set, Hamiltonian Cycle, remain so on $\mathcal{T}_{4}$. However, Min Coloring and $k$-Coloring are easy on both families because they are 2-colorable and 3-colorable, respectively.
We show that Min Coloring is $\mathsf{NP}$-hard on the class $\mathcal{T}_{3}$ of graphs of twin-width at most 3. This is the first hardness result on $\mathcal{T}_{3}$ for a problem that is easy on cographs (twin-width 0), on trees (whose twin-width is at most 2), and on unit circular-arc graphs (whose twin-width is at most 3). We also show that for every $\geqslant 3, kk$-Coloring is $\mathsf{NP}$-hard on $\mathcal{T}_{4}$. We finally make two observations: (1) there are currently very few problems known to be in $\mathsf{P}$ on $\mathcal{T}_{d}$ (graphs of twin-width at most $d$) and $\mathsf{NP}$-hard on $\mathcal{T}_{d+1}$ for some nonnegative integer $d$, and (2) unlike $\mathcal{T}_{4}$, which contains every graph as an induced minor, the class $\mathcal{T}_{3}$ excludes a fixed planar graph as an induced minor; thus it may be viewed as a special case (or potential counterexample) for conjectures about classes excluding a (planar) induced minor. These observations are accompanied by several open questions.
category:
[…]
1 Introduction
The graph parameter twin-width was introduced in 2020 [15]. The family of graph classes of bounded twin-width is broad and diverse: it includes classes of bounded clique-width, $d$-dimensional grids, classes excluding a fixed minor [15], some cubic expanders [11]. Moreover, there are algorithms that work on any class of effectively [^1] A class $\mathcal{C}$ has effectively bounded twin-width if there is an integer $d$ and a polynomial-time algorithm that outputs a $d$-sequence (see Section 2 for all relevant definitions) of any input graph from $\mathcal{C}$. bounded twin-width: a fixed-parameter tractable first-order model checking algorithm [15], single-exponential parameterized algorithms [12], improved approximation algorithms [12, 6], fast shortest-path algorithms [12, 4], etc. Most $\mathsf{NP}$-hard graph problems remain intractable on graphs of twin-width at most 4. This paper tackles the hardness of coloring graphs of low twin-width.
For any nonnegative integer $d$, we denote by $\mathcal{T}_{d}$ the class of graphs of twin-width at most $d$. Finite subgraphs of the infinite planar grid [15] and $(\geqslant 2\log n)$-subdivisions of $n$-vertex graphs [5] are in $\mathcal{T}_{4}$. Problems such as Max Independent Set, Vertex Cover, Max Clique, Feedback Vertex Set, Dominating Set, and Max Induced Matching are $\mathsf{NP}$-hard on the latter family, while Hamiltonian Path and Hamiltonian Cycle are $\mathsf{NP}$-hard on the former. Therefore all these problems are $\mathsf{NP}$-hard on $\mathcal{T}_{4}$. Notably, Min Coloring is easy on subgraphs of the grid, since they are bipartite. It is also easy on strict subdivisions (i.e., when every edge is subdivided at least once), since such graphs are always 3-colorable and bipartiteness can be checked in linear time. We show that Min Coloring is already $\mathsf{NP}$-hard on $\mathcal{T}_{3}$. This is the first problem that is shown $\mathsf{NP}$-hard on $\mathcal{T}_{3}$ while being in $\mathsf{P}$ on classes known to have twin-width at most 3 (cographs, trees, unit circular-arc graphs).
Theorem 1.1.
Min Coloring is $\mathsf{NP}$-hard and, unless the ETH fails, requires $2^{\Omega(\sqrt{n})}$ time on $n$-vertex graphs of twin-width at most 3 (even if a 3-sequence is provided).
In the previous theorem, the ETH (for Exponential-Time Hypothesis) [29] asserts that there exists a constant $\lambda>1$ such that $n$-variable 3-Sat cannot be solved in time $O(\lambda^{n})$. The reduction requires the number of colors to grow with $n$.
We also show that 3-Coloring is $\mathsf{NP}$-hard on $\mathcal{T}_{4}$. Note that planar graphs have twin-width at least 7 [34] (and at most 8 [28]). So we cannot just invoke the hardness of 3-Coloring in planar graphs. We tune long subdivisions to turn them into hard 3-Coloring instances.
Theorem 1.2.
3-Coloring is $\mathsf{NP}$-hard on graphs of twin-width at most 4.
This again holds even if a 4-sequence is provided and directly implies that $k$-Coloring is also hard on $\mathcal{T}_{4}$, for any fixed $k\geqslant 3$.
On the algorithmic side, most graph problems are in $\mathsf{P}$ on $\mathcal{T}_{1}$ since this class has bounded clique-width [13] and is a subclass of permutation graphs [3]. Although 1-sequences can be computed in polynomial time in $\mathcal{T}_{1}$ [14], and even in linear time [3], the corresponding algorithms do not require a sequence to be provided as part of the input. Max Independent Set (and consequently, Vertex Cover and Max Clique) can be solved in polynomial time in $\mathcal{T}_{2}$ if 2-sequences are provided as part of the input [12].
While $\mathcal{T}_{1}$ can be recognized in polynomial time, it is $\mathsf{NP}$-hard to decide if a graph (of $\mathcal{T}_{5}$) is in $\mathcal{T}_{4}$. The complexity of the recognition of $\mathcal{T}_{2}$ and of $\mathcal{T}_{3}$ is open.
$\mathcal{T}_{1}$ $\mathcal{T}_{2}$ $\mathcal{T}_{3}$ $\mathcal{T}_{4}$
Min Coloring $\mathsf{P}$ [3, 22] ? $\mathsf{NP}$-c (Theorem˜1.1)
3-Coloring $\mathsf{P}$ [13, 16] ? ? $\mathsf{NP}$-c (Theorem˜1.2)
$k$-Coloring, $k\geqslant 3$ $\mathsf{P}$ [13, 16] ? ? $\mathsf{NP}$-c (Corollary˜4.2)
Max Independent Set $\mathsf{P}$ [13, 16] $\mathsf{P}$ [12] ? $\mathsf{NP}$-c [5]
Vertex Cover $\mathsf{P}$ [13, 16] $\mathsf{P}$ [12] ? $\mathsf{NP}$-c [5]
Max Clique $\mathsf{P}$ [13, 16] $\mathsf{P}$ [12] ? $\mathsf{NP}$-c [5]
Feedback Vertex Set $\mathsf{P}$ [13, 16] ? ? $\mathsf{NP}$-c [5]
Dominating Set $\mathsf{P}$ [13, 16] ? ? $\mathsf{NP}$-c [5]
Max Induced Matching $\mathsf{P}$ [13, 16] ? ? $\mathsf{NP}$-c [5]
Hamiltonian Path $\mathsf{P}$ [3, 20] ? ? $\mathsf{NP}$-c [15, 31]
Hamiltonian Cycle $\mathsf{P}$ [3, 21] ? ? $\mathsf{NP}$-c [15, 31]
Recognition $\mathsf{P}$ [14, 3] ? ? $\mathsf{NP}$-c [5]
Table 1: Complexity of some of the main $\mathsf{NP}$-complete graph problems in $\mathcal{T}_{d}$ for $d=1,2,3,4$. The $\mathsf{P}$ in gray in the $\mathcal{T}_{2}$ column means that 2-sequences are required by the known algorithm.
Table˜1 suggests the task of closing these gaps.
Question 1.3.
Establish, for a problem $\Pi$ of Table˜1, a nonnegative integer $d$ such that $\Pi$ on $\mathcal{T}_{d}$ is in $\mathsf{P}$ and $\Pi$ on $\mathcal{T}_{d+1}$ is $\mathsf{NP}$-hard.
To our knowledge, the only problems for which Question˜1.3 is settled are Firefighter and Restricted Vertex Multicut, which are in $\mathsf{P}$ on $\mathcal{T}_{1}$ (as they are polynomial-time solvable on permutation graphs [25, 35]) and $\mathsf{NP}$-hard on $\mathcal{T}_{2}$ (as they are hard on trees [24, 17]).
As the current paper is on graph coloring, we ask the following question.
Question 1.4.
Is Min Coloring on $\mathcal{T}_{2}$ in $\mathsf{P}$?
As we mentioned, Min Coloring is the first problem to be shown $\mathsf{NP}$-hard on $\mathcal{T}_{3}$ while being tractable on classes known to have twin-width at most 3: cographs (which coincide with $\mathcal{T}_{0}$ [15]), trees (which are in $\mathcal{T}_{2}$ [15]), and unit circular-arc graphs (which are in $\mathcal{T}_{3}$ [9, below Theorem 5.5]). Without this requirement, other examples would include Achromatic Number, which is $\mathsf{NP}$-hard on cographs [8], Bandwidth, which is $\mathsf{NP}$-hard on trees [26], and Partial Representation Extension, which is $\mathsf{NP}$-hard on unit circular-arc graphs [23]. Theorem˜1.1 thus involved a novel encoding that incurred a quadratic blow-up. We wonder whether this blow-up can be avoided.
Question 1.5.
Can Min Coloring be solved in time $2^{O(\sqrt{n})}$ on $n$-vertex graphs of $\mathcal{T}_{3}$?
The class $\mathcal{T}_{1}$ is now well understood [3]. In particular, it is a subclass of permutation graphs with bounded clique-width. We believe that the classes $\mathcal{T}_{2}$ and $\mathcal{T}_{3}$ have some still hidden structure. While the membership problem in $\mathcal{T}_{4}$ is $\mathsf{NP}$-hard [5], that of $\mathcal{T}_{2}$ and of $\mathcal{T}_{3}$ are open. It was proven that weakly sparse subclasses (i.e., excluding a biclique $K_{t,t}$ as a subgraph) of $\mathcal{T}_{2}$ have bounded treewidth, whereas this is not the case for $\mathcal{T}_{3}$ [7].
However, bounded-degree subclasses of $\mathcal{T}_{3}$ have bounded treewidth. This is because bounded-degree graphs of large treewidth admit subdivisions of large walls or their line graphs as induced subgraphs [33] and those graphs have twin-width (exactly) 4 [2]. Furthermore, the latter result combined with [1, 10] (extending [33]) implies that every subclass of $\mathcal{T}_{3}$ without large subdivided cliques as subgraphs has bounded treewidth. As 3-Coloring is in $\mathsf{P}$ on any class of bounded treewidth, if 3-Coloring is $\mathsf{NP}$-hard on $\mathcal{T}_{3}$, then the hard instances must contain arbitrarily large clique subdivisions as subgraphs. The hard 3-Coloring instances that the proof of Theorem˜1.2 builds do not: they have a single vertex of degree more than 4.
Question 1.6.
Is 3-Coloring on $\mathcal{T}_{3}$ in $\mathsf{P}$?
The same question holds for Max Independent Set.
Question 1.7.
Is Max Independent Set on $\mathcal{T}_{3}$ in $\mathsf{P}$?
Another consequence of the above result of [2] is that $\mathcal{T}_{3}$ excludes a fixed planar graph as an induced minor. Thus Questions˜1.7 and 1.6 are special cases of the same questions on any class excluding a planar induced minor, which was previously raised for Max Independent Set [19]. The class $\mathcal{T}_{3}$ is a good candidate for a negative answer to the latter question (on the pessimistic side), or Question˜1.7 could serve as a preliminary step in positively answering it (on the optimistic side). There are other questions and conjectures on classes excluding a planar induced minor. They can be revisited on the particular class $\mathcal{T}_{3}$, like for instance the following conjecture appearing in [27].
Question 1.8.
Is there a universal constant $c$ such that every graph of $\mathcal{T}_{3}$ admits a balanced separator included in the neighborhood of at most $c$ vertices?
We recall the relevant definitions and notation in Section˜2, show Theorem˜1.1 in Section˜3, and Theorem˜1.2 in Section˜4.
2 Preliminaries
For two integers $i$ and $j$, we denote by $[i,j]$ the set of integers that are at least $i$ and at most $j$. For every integer $i$, $[i]$ is a shorthand for $[1,i]$.
2.1 Standard graph-theoretic definitions and notation
We denote by $V(G)$ and $E(G)$ the set of vertices and edges of a graph $G$, respectively. For $S\subseteq V(G)$, the subgraph of $G$ induced by $S$, denoted $G[S]$, is obtained by removing from $G$ all the vertices that are not in $S$. Then GSG-S is a shorthand for $G[V(G)\setminus S]$. We denote by $N_{G}(v)$ the set of neighbors of $v$ in $G$. A subdivision of a graph $G$ is any graph $H$ obtained from $G$ by replacing edges $e$ of $G$ by paths with at least one edge whose extremities are the endpoints of $e$. A $(\geqslant s)$-subdivision is a subdivision where every edge is replaced by a path with at least $s$ internal vertices.
In this paper, a coloring of a graph is implicitly assumed to be a proper vertex-coloring.
2.2 Trigraphs, partition sequences, and twin-width
[…]
Given a (tri)graph $G$ and a partition $\mathcal{P}$ of $V(G)$, the quotient trigraph $G/\mathcal{P}$ is the trigraph with vertex set $\mathcal{P}$, where $PP^{\prime}$ is a black edge if these two parts are fully adjacent via black edges (i.e., for every $u\in P$ and every $v\in P^{\prime}$, $uv\in E(G)$), and a red edge if there is $u\in P$ and $v\in P^{\prime}$ such that $uv\in R(G)$ or $u_{1},u_{2}\in P$ and $v_{1},v_{2}\in P^{\prime}$ such that $u_{1}v_{1}\in E(G)$ and $u_{2}v_{2}\notin E(G)$.
A partition sequence of an $n$-vertex (tri)graph $G$ is a sequence $\mathcal{P}_{n},\mathcal{P}_{n-1},\ldots,\mathcal{P}_{1}$ of partitions of $V(G)$ such that $\mathcal{P}_{n}=\{\{v\}:v\in V(G)\}$, $\mathcal{P}_{1}=\{V(G)\}$, and for every $i\in[n-1]$, $\mathcal{P}_{i}$ is obtained from $\mathcal{P}_{i+1}$ by merging $P,P^{\prime}\in\mathcal{P}_{i+1}$ into $P\cup P^{\prime}$. A $d$-sequence of $G$ is a partition sequence $\mathcal{P}_{n},\ldots,\mathcal{P}_{1}$ such that for every $i\in[n]$, the maximum red degree of $G/\mathcal{P}_{i}$ is at most $d$. The twin-width of a (tri)graph is the least integer $d$ such that it admits a $d$-sequence.
A partition sequence is called partial when we relax the condition that the last partition has a single part. We say that a trigraph is fully red if it does not have any black edges. We will use the simple fact that turning some (or all) black edges red cannot decrease the twin-width of a trigraph. Therefore, when showing twin-width upper bounds, we may sometimes assume that all the edges (black and red) are in fact red, if this simplifies a later argument.
3 Hardness of Coloring Graphs of Twin-Width 3
As a decision problem, Min Coloring takes as input a graph $G$ and an integer $k$ and asks whether the chromatic number of $G$, $\chi(G)$, is at most $k$. As a function problem, one is given the mere graph $G$ and has to output a (proper) $\chi(G)$-coloring of $G$. We show that Min Coloring is $\mathsf{NP}$-hard on $\mathcal{T}_{3}$, even in its decision form and when a 3-sequence of $G$ is given as input. As Min Coloring is polynomial-time solvable on cographs, on trees, and on unit circular-arc graphs (all of which have twin-width at most 3), and planar graphs are not included in $\mathcal{T}_{3}$, this requires finding a novel kind of encoding of rich structures (here, 3-Sat instances) onto graphs of twin-width at most 3.
Theorem 1.1.
Min Coloring is $\mathsf{NP}$-hard and, unless the ETH fails, requires $2^{\Omega(\sqrt{N})}$ time on $N$-vertex graphs of twin-width at most 3 (even if a 3-sequence is provided).
Proof 3.1.
We reduce from 3-Sat, which is $\mathsf{NP}$-complete [32], and unless the ETH fails, its $n$-variable $m$-clause instances cannot be solved in time $2^{o(n+m)}$ by the so-called Sparsification Lemma [30]; it is indeed shown that, under the ETH, $n$-variable 3-Sat cannot be solved in time $2^{o(n)}$ even on instances with $O(n)$ clauses. Let $x_{1},\ldots,x_{n}$ be the variables of a 3-Sat instance $\varphi$, and $c_{1},\ldots,c_{m}$ be its clauses.
We describe the construction of a graph $G:=G(\varphi)$ with chromatic number (at most) $2n$ if and only if $\varphi$ is satisfiable. We start with two paths $P,P^{\prime}$ each on $2np$ vertices with $p:=2n+m$, which we partition evenly into $p$ subpaths on $2n$ vertices. Let $A_{1},A_{2},\ldots,A_{p}$ (resp. $B_{1},B_{2},\ldots,B_{p}$) be the vertex sets of the subpaths of $P$ (resp. $P^{\prime}$) from left to right. We set $A:=\bigcup_{i\in[p]}A_{i}=V(P)$ and $B:=\bigcup_{i\in[p]}B_{i}=V(P^{\prime})$. We denote by $a_{i,1},a_{i,2},\ldots,a_{i,2n}$ the $2n$ vertices of the subpath $P[A_{i}]$ from left to right. Similarly we denote by $b_{i,1},b_{i,2},\ldots,b_{i,2n}$ the $2n$ vertices of the subpath $P^{\prime}[B_{i}]$ from left to right. We now define $G[A]$ (resp. $G[B]$) as $P^{\leqslant 2n-1}$ (resp. $P^{\prime\leqslant 2n-1}$), that is, for every $u,v\in A$ (resp. $u,v\in B$), $uv\in E(G)$ whenever $u$ and $v$ are at distance at most 2n12n-1 in $P$ (resp. in $P^{\prime}$). In particular, each $A_{i}$ (and each $B_{i}$) is a clique of size $2n$, and for all $i,i^{\prime}\in[p]$ with $i<i^{\prime}$, it holds that $a_{i,j}$ and $a_{i^{\prime},j^{\prime}}$ are adjacent if and only if $i^{\prime}=i+1$ and $j^{\prime}<j$. There is no edge between $A$ and $B$.
For every $i\in[p]$, we add one vertex $v_{i}$ adjacent to a subset of $A_{i}\cup B_{i}$, as follows.
  • For every $i\in[n]$, $N_{G}(v_{2i-1})=(A_{2i-1}\cup B_{2i-1})\setminus\{a_{2i-1,2i-1},a_{2i-1,2i},b_{2i-1,2i-1}\}$ and $N_{G}(v_{2i})=(A_{2i}\cup B_{2i})\setminus\{a_{2i,2i-1},a_{2i,2i},b_{2i,2i}\}$.
  • For every $i\in[2n+1,2n+m]$, say $c_{i-2n}$ is the clause $s_{1}x_{j_{1}}\lor s_{2}x_{j_{2}}\lor s_{3}x_{j_{3}}$ (repeat a literal if it had only two), where $s_{1},s_{2},s_{3}$ are signs in $\{\neg,\varepsilon\}$ ($\varepsilon$ is the positive sign). Then, $N_{G}(v_{i})=(B_{i}\setminus\{b_{i,2j_{1}},b_{i,2j_{2}},b_{i,2j_{3}}\})\cup\{a_{i,2j_{1}-f(s_{1})},a_{i,2j_{2}-f(s_{2})},a_{i,2j_{3}-f(s_{3})}\}$ with $f(\neg)=0$ and $f(\varepsilon)=1$.
This finishes the construction of $G$; see Figure˜1 for an illustration. Note that $G$ has $N:=(4n+1)p=O(n(n+m))$ vertices, and that it can be constructed in polynomial time from $\varphi$. Thus, provided the reduction is correct, which we next prove, a $2^{o(\sqrt{N})}$-time algorithm for Min Coloring on the produced instances would imply a $2^{o(n+m)}$-time algorithm for $n$-variable $m$-clause 3-Sat, hence refute the ETH.
$A_{1}$ $B_{1}$$v_{1}$ $A_{2}$ $B_{2}$$v_{2}$ $A_{3}$ $B_{3}$$v_{3}$ $A_{4}$ $B_{4}$$v_{4}$ $A_{5}$ $B_{5}$$v_{5}$ $A_{6}$ $B_{6}$$v_{6}$ $A_{7}$ $B_{7}$$v_{7}$ $A_{8}$ $B_{8}$$v_{8}$
Figure 1: The graph $G(\varphi)$ for $\varphi$ consisting of two clauses $c_{1}=x_{1}\lor\neg x_{2}\lor x_{3}$ and $c_{2}=\neg x_{1}\lor x_{2}\lor\neg x_{3}$. The variable gadgets and clause gadgets are delimited by the vertical dashed line. Not to clutter the picture, we have not drawn the edges in $G[A\cup B]$.
If $\bm{\varphi}$ is satisfiable, then $\bm{G}$ is $\bm{2n}$-colorable. Let $\mathcal{A}$ be a satisfying assignment for $\varphi$. We then describe a (proper) $2n$-coloring $c$ of $G$.
  • For every $i\in[p]$ and every $j\in[2n]$, we set $c(b_{i,j})=j$.
  • For every $i\in[p]$ and every $j\in[n]$, we set $c(a_{i,2j-1})=2j-1$ and $c(a_{i,2j})=2j$ if $\mathcal{A}$ assigns $x_{j}$ to true, or $c(a_{i,2j})=2j-1$ and $c(a_{i,2j-1})=2j$, otherwise.
As the $2n$ colors repeat with period $2n$ along each path of $\{P,P^{\prime}\}$, no edge of $G[A\cup B]$ is monochromatic. We now simply need to argue that for every vertex $v_{i}$, there is at least one color $c^{\prime}$ of $[2n]$ not present in $N_{G}(v_{i})$; and set $c(v_{i})=c^{\prime}$.
For every $i\in[2n]$, we can set $c(v_{i})=i$, as this color does not appear in $N_{G}(v_{i})$. For every $i\in[2n+1,2n+m]$, with $c_{i-2n}=s_{1}x_{j_{1}}\lor s_{2}x_{j_{2}}\lor s_{3}x_{j_{3}}$ and $s_{1},s_{2},s_{3}\in\{\neg,\varepsilon\}$, there is an $h\in\{1,2,3\}$ such that $\mathcal{A}$ satisfies $s_{h}x_{j_{h}}$. Then we can set $c(v_{i})=2j_{h}$. Indeed, if $\mathcal{A}$ sets $x_{j_{h}}$ to true, then $f(s_{h})=1$ and $a_{i,2j_{h}-1}$ is colored $2j_{h}-1$, whereas if $\mathcal{A}$ sets $x_{j_{h}}$ to false, then $f(s_{h})=0$ and $a_{i,2j_{h}}$ is colored $2j_{h}-1$; thus, in both cases, color $2j_{h}$ is still available for $v_{i}$. See Figure˜2 for an example.
$A_{1}$ $B_{1}$$v_{1}$ $A_{2}$ $B_{2}$$v_{2}$ $A_{3}$ $B_{3}$$v_{3}$ $A_{4}$ $B_{4}$$v_{4}$ $A_{5}$ $B_{5}$$v_{5}$ $A_{6}$ $B_{6}$$v_{6}$ $A_{7}$ $B_{7}$$v_{7}$ $A_{8}$ $B_{8}$$v_{8}$
Figure 2: A (proper) $2n$-coloring of the graph $G(\varphi)$ of Figure˜1 corresponding to the satisfying assignment $\mathcal{A}=\{x_{1}\mapsto\text{false},x_{2}\mapsto\text{true},x_{3}\mapsto\text{true}\}$. The coloring of $v_{7}$ (resp. $v_{8}$) witnesses that $c_{1}$ (resp. $c_{2}$) is satisfied by literal $x_{3}$ (resp. $\neg x_{1}$).
If $\bm{G}$ is $\bm{2n}$-colorable, then $\bm{\varphi}$ is satisfiable. Fix a (proper) coloring $c\colon V(G)\to[2n]$ of $G$, with $c(S):=\{c(v):v\in S\}$ for any $S\subseteq V(G)$. As $B_{1}$ is a clique, we can further assume that $c(b_{1,j})=j$ for every $j\in[2n]$. By a straightforward induction, for every $i\in[p-1]$ and $j\in[2n]$, $c(a_{i,j})=c(a_{i+1,j})$ and $c(b_{i,j})=c(b_{i+1,j})=j$, and for every $i\in[p]$, $c(A_{i})=[2n]=c(B_{i})$. Indeed both $A_{1}$ and $B_{1}$ are $2n$-vertex cliques, thus $c(A_{1})=c(B_{1})=[2n]$, and when $2n$-coloring $P^{\leqslant 2n-1}$ (resp. $P^{\prime\leqslant 2n-1}$) from left to right, the only available color for $a_{i+1,j}$ (resp. $b_{i+1,j}$) is $c(a_{i,j})$ (resp. $c(b_{i,j})$).
We first show that, for every $i\in[p]$ and $j\in[n]$, $\{c(a_{i,2j-1}),c(a_{i,2j})\}=\{2j-1,2j\}$. Assume for contradiction that there is an $h\in\{c(a_{i,2j-1}),c(a_{i,2j})\}\setminus\{2j-1,2j\}$. By the previous paragraph, observe that $h$ does not depend on $i$. This implies that $2j-1\notin\{c(a_{i,2j-1}),c(a_{i,2j})\}$ or $2j\notin\{c(a_{i,2j-1}),c(a_{i,2j})\}$. In the former case, $v_{2j-1}$ would need a $(2n+1)$-st color, whereas in the latter case, the same would happen to $v_{2j}$.
Let $\mathcal{A}$ be the truth assignment that sets, for every $j\in[n]$, $x_{j}$ to true if $c(a_{i,2j-1})=2j-1$ and $c(a_{i,2j})=2j$, and to false, if $c(a_{i,2j-1})=2j$ and $c(a_{i,2j})=2j-1$ (again, note that this does not depend on $i$). Let us show that $\mathcal{A}$ satisfies all the clauses of $\varphi$. For each $i\in[2n+1,2n+m]$, say $c_{i-2n}=s_{1}x_{j_{1}}\lor s_{2}x_{j_{2}}\lor s_{3}x_{j_{3}}$ with $s_{1},s_{2},s_{3}\in\{\neg,\varepsilon\}$. Since $v_{i}$ is adjacent to $B_{i}\setminus\{b_{i,2j_{1}},b_{i,2j_{2}},b_{i,2j_{3}}\}$, it holds that $c(v_{i})\in\{2j_{1},2j_{2},2j_{3}\}$; say, $c(v_{i})=2j_{h}$ with $h\in\{1,2,3\}$. This implies that $c(a_{i,2j_{h}-f(s_{h})})\neq 2j_{h}$, thus $c(a_{i,2j_{h}-f(s_{h})})=2j_{h}-1$. Therefore, $s_{h}=\varepsilon$ and $\mathcal{A}$ sets $x_{j_{h}}$ to true, or $s_{h}=\neg$ and $\mathcal{A}$ sets $x_{j_{h}}$ to false. In both cases, the literal $s_{h}x_{j_{h}}$ of $c_{i-2n}$ is satisfied by $\mathcal{A}$.
$\bm{G}$ admits a 3-sequence. We describe a partition sequence for $G$ of width at most 3. The sequence has two stages. In the first stage, each $A_{i}$ and each $B_{i}$ is merged into a single part. Throughout this stage, we denote by $P_{i}$ (resp. $P^{\prime}_{i}$) the current part containing $a_{i,1}$ (resp. $b_{i,1}$). For $j$ going from 2 to $2n$, for $i$ going from 1 to $p$, merge $P_{i}$ and singleton $\{a_{i,j}\}$, and merge $P^{\prime}_{i}$ and singleton $\{b_{i,j}\}$.
Let us argue that this partial sequence has width at most 3. At any point, the singleton parts $\{a_{i,j+1}\},\ldots,\{a_{i,2n}\}$ in between $P_{i}=\{a_{i,1},\ldots,a_{i,j}\}$ and $P_{i+1}$ are all fully adjacent to $P_{i}\cup P_{i+1}$, some of these singletons are adjacent to $v_{i}$, and they have no other adjacencies. Thus these parts have red degree 0. Note that the singleton parts $\{a_{p,j+1}\},\ldots,\{a_{p,2n}\}$ do not have incident red edges either. Each singleton part $\{v_{i}\}$ has red degree at most 2, without any red neighbor outside $\{P_{i},P^{\prime}_{i}\}$. Finally, at any point, each $P_{i}$ has at most three red neighbors: $P_{i-1}$ (if it exists), $v_{i}$, and $P_{i+1}$ (if it exists). Indeed, all the other parts are either fully nonadjacent to $P_{i}$ or have red degree 0. Symmetrically, $P^{\prime}_{i}$ has always at most three red neighbors.
After the first stage is completed, the resulting trigraph is fully red (we may well assume that $n\geqslant 2$), has $3p$ vertices, and consists of two $p$-vertex paths whose $i$th vertices have an additional shared neighbor of degree 2, for every $i\in[p]$. In the second stage, we finish the 3-sequence, as follows. Let us call $Q$ the part containing $v_{1}$. For $i$ going from 2 to $p$, we merge $P_{1}$ and $P_{i}$, and keep denoting the resulting part by $P_{1}$, we merge $P^{\prime}_{1}$ and $P^{\prime}_{i}$, and denote the resulting part by $P^{\prime}_{1}$, and then we merge $Q$ and $\{v_{i}\}$. Finally, only three parts remain that we merge to a single part in any way. One can observe that at any time of the second stage when at least three parts are left, every part $P_{i}$ or $P^{\prime}_{i}$ (including $P_{1},P^{\prime}_{1}$) has at most three red neighbors, and $Q$, as well as every part $\{v_{i}\}$, has two red neighbors. So this finishes the description of a 3-sequence of $G(\varphi)$.
As this sequence can be computed in polynomial time for any formula $\varphi$, the hardness of Min Coloring on $\mathcal{T}_{3}$ further holds when a 3-sequence is provided as part of the input.
4 Hardness of 3-Coloring Graphs of Twin-Width 4
The hardness proof of the previous section requires an unbounded number of colors. In this section we show that deciding if a graph of twin-width at most 4 is 3-colorable is $\mathsf{NP}$-hard. We also get a similar ETH lower bound, and readily extend this hardness result to $k$-Coloring.
Theorem 1.2.
3-Coloring is $\mathsf{NP}$-hard and, unless the ETH fails, requires $2^{\Omega(n/\log n)}$ time on $n$-vertex graphs of twin-width at most 4 (even if a 4-sequence is provided).
Proof 4.1.
It is known that 3-Coloring is $\mathsf{NP}$-hard on $n$-vertex graphs of maximum degree at most 4, and requires $2^{\Omega(n)}$ time, unless the ETH fails (see for instance [18, Lemma 1]).
Let $G$ be an $n$-vertex graph of degree at most 4. We build an $O(n\log n)$-vertex graph $G^{\prime}$ of twin-width at most 4 such that $G^{\prime}$ is 3-colorable if and only if $G$ is 3-colorable. Let $q:=2\lceil\log n\rceil$. Graph $G^{\prime}$ is obtained from $G$ by subdividing each edge $e\in E(G)$ $q$ times, thereby creating vertices $v_{e,1},\ldots,v_{e,q}$, and then adding one true twin (i.e., an adjacent vertex with the same closed neighborhood) $v^{\prime}_{e,1},v^{\prime}_{e,3},v^{\prime}_{e,5},\ldots,v^{\prime}_{e,q-1}$ to $v_{e,1},v_{e,3},v_{e,5},\ldots,v_{e,q-1}$, respectively (that is, every other created vertex gets a true twin). The “orientation” of $e$ to fully specify this process is irrelevant, so we leave it undefined. (To get a unique construction, one can totally order $V(G)$, and always “orient” the edge from its smaller endpoint to its larger endpoint.)
The way $G^{\prime}$ is built, it can be observed that there is a partial 0-sequence from $G^{\prime}$ to the $q$-subdivision of $G$. Since every $(\geqslant 2\log n)$-subdivision of an $n$-vertex graph has twin-width at most 4 [5], $G^{\prime}$ has twin-width at most 4. The number of vertices of $G^{\prime}$ is $O(n+|E(G)|\log n)=O(n\log n)$, and its construction can be done in $O(n\log n)$ time.
Given an original vertex $u$ (i.e., $u\in V(G)$) of $G^{\prime}$ adjacent to some $v_{e,1}$, and a color $i\in[3]$, the 3-coloring propagation of $u$ along $e$ is the only (proper) 3-coloring of
$G[\{u,v_{e,1},v^{\prime}_{e,1},v_{e,2},v_{e,3},v^{\prime}_{e,3},\ldots,v_{e,q-1},v^{\prime}_{e,q-1},v_{e,q}\}]$
that gives color $i$ to $u$. Note that this 3-coloring also gives color $i$ to $v_{e,q}$.
If $G$ is 3-colorable, we start by copying a 3-coloring of $G$ at the original vertices of $G^{\prime}$. We then make the 3-coloring propagation for each pair $u\in V(G),v_{e,1}$ with $uv_{e,1}\in E(G^{\prime})$. By the previous remark, this defines a (proper) 3-coloring of $G^{\prime}$. Conversely, the projection of any 3-coloring of $G^{\prime}$ onto $V(G)$ properly colors $G$.
Corollary 4.2.
For every $\geqslant 3, kk$-Coloring is $\mathsf{NP}$-hard and, unless the ETH fails, requires $2^{\Omega(n/\log n)}$ time on $n$-vertex graphs of twin-width at most 4.
Proof 4.3.
Adding k3k-3 universal vertices to the graph $G^{\prime}$ built in the proof of Theorem˜1.2 does not change its twin-width and makes the resulting graph $k$-colorable if and only if $G^{\prime}$ is 3-colorable.